Consider the following model of national income determination
C = 300 + 0.75 (Y-T)
T = 100
I = 475
G = 150

i) List the entire exogenous and endogenous variable. (2 marks)

ii) Solve for the equilibrium value for all the endogenous variables. (2 marks)

iii) Suppose government expenditure increase by 50 find the new equilibrium values of the endogenous variables. Assume the economy of Kenya is described by the following information
Y = C + I + G (x - M) x = 20
C=20+0.8Yd M = 4 + 0.3Y
T = 30 Yd=Y-T
G = 22 I = 30 (8 marks)
1
Expert's answer
2018-01-23T08:36:08-0500
i) C - endogenous variable
Y - exogenous variable
ii) Y = 300 + 0,75(Y-100)+475+150=3400 2775 , C = 2775
iii) C = 2812,5 , Y = 3450
Y = C + I + G+ (x-M)=128 for Keniya